in a ideal gas if the pressure is multilplied by 2 and the temperature remains the same, what is the new volume? is it halved or the volume doesnt change because an ideal gas depends on the temperature
Answer #1

given data

let

Initial pressure P1= P

INITIAL volume = V1

initial TEMPERATURE =T1 =T

final pressure P2= 2P

Final temperature T2 =T

final volume V2= ?

WE KNOW THAT

( P1V1/T1 ) =( P2V2/T2)

PLUGGING THE VALUES

PV1/T = 2PV2/T

V2= (1/2)V1

ANS IS VOLUME IS HALVED

It means that for a gas at a constant temperature, pressure × volume is also constant. So increasing pressure from pressure1 to pressure2 means that volume1 will change to volume2, providing the temperature remains constant

ans is HALVED

IF YOU FIND THIS HELPFUL PLEASE UP VOTE